7. The payback period Aa Aa E The payback method helps firms establish and identify a maximum acceptable payback period thatThe conventional payback period ignores the time value of money, and this concerns Cute Camels CFO. He has now asked you to
Answer #1

E 1 2 Expected cash flow 3 Cumulative cash flow 4 5 Conventional payback period Year 0 (6,000,000) (6,000,000) c ID Year 1 YeYear 0 Year 1 Year 2 Year 3 -6000000 =B2 2400000 =B3+C2 5100000 =C3+D2 2100000 =D3+E2 2 Expected cash flow 3 Cumulative cash